b) Find the power delivered to an element at t = 3 ms if the current entering its positive terminal is i = 5cos60πt A and the voltage is

(i) v = 3i

(ii) v = 3di/dt         

1
Expert's answer
2022-03-24T16:18:49-0400

"P=iv,"

a)

"v=15\cos60\pi t,"

"P=75\cos^260\pi t,"

"P(3)=75~W,"

b)

"v=-900\pi \sin 60\pi t,"

"P=-2250\pi \sin 120\pi t,"

"P(3)=0."